# Parameter estimation for Gibbs distributions

### v9

@@ 6.2 Converting the preschedule into a covering schedule
From Proposition 38, we have ${\tt InvWt}(\mathcal{I})=\sum_{i=0}^{t}\frac{1}{e^{-\nu}w_{i}}\leq\frac{2e^{\nu}(n+1)}{\phi}$. Similarly, the algorithm cost from Line 6 is $\sum_{i=0}^{t}O(w_{i}\log\tfrac{t}{\gamma})\leq O(\frac{2(n+1)}{\phi}\cdot\log\tfrac{n}{\gamma})\leq O(n\rho\log\tfrac{n}{\gamma})$, thus showing part (d).
To show $\mathcal{I}$ is a covering schedule, we need to show that $k_{1}<\dots<k_{t}$ strictly and $\beta_{0}=\beta_{\min}<\dots<\beta_{t}=\beta_{\max}$ strictly. The condition $\beta_{0}=\beta_{\min},\beta_{t}=\beta_{\max}$ follows from (I1) while the bound $\beta_{i}<\beta_{i+1}$ follows from the loop at Line 1. To show $k_{i}<k_{i+1}$ for $i\in\{1,\dots,t-1\}$, observe that from (I0) we have $\sigma_{i-1}^{+}\geq\sigma_{i}^{-}$ and $\sigma_{i}^{+}\geq\sigma_{i+1}^{-}$. Due to the loop at Line 3, we must have $k_{i+1}\geq\sigma_{i+1}^{-}>\sigma_{i-1}^{+}\geq k_{i}$ (else segment $\theta_{i}$ would have been discarded).
Now suppose that Line 6 well-estimates each $\sigma_{i}^{+},\sigma_{i}^{-},\sigma_{i+1}^{-},\sigma_{i+1}^{+}$; by specification of the parameters this has probability at least $1-\gamma$. We claim that the algorithm outputs either a proper covering schedule or ERROR, and that the latter case only holds if $\mathcal{J}^{\prime}$, and hence $\mathcal{J}$, is improper.
First, if the algorithm reaches Line 11, then $\hat{\mu}_{\beta_{i}}(k_{i})\geq e^{-\nu/2}w_{i}$ and $\hat{\mu}_{\beta_{i-1}}(k_{i})\geq e^{-\nu/2}w_{i-1}$ for each $i$. Since Line 6 well-estimates $k_{i}$, this implies $\mu_{\beta_{i}}(k_{i})\geq e^{-\nu}w_{i}$ and $\mu_{\beta_{i-1}}(k_{i})\geq e^{-\nu}w_{i-1}$. So $\mathcal{I}$ is proper.
Next, suppose $\mathcal{J}^{\prime}$ is $\phi$-proper but the algorithm outputs ERROR at an iteration $i$. By definition of $\phi$-properness, we have $\mu_{\beta_{i-1}}(\sigma_{i-1}^{+})\geq w_{i-1}$ and $\mu_{\beta_{i}}(\sigma_{i}^{-})\geq w_{i}$. Since Line 6 well-estimates $\sigma_{i-1}^{+}$ and $\sigma_{i}^{+}$, this implies $\hat{\mu}_{\beta_{i-1}}(\sigma_{i-1}^{+})\geq e^{-\nu/2}w_{i-1}$ and $\hat{\mu}_{\beta_{i}}(\sigma_{i}^{-})\geq e^{-\nu/2}w_{i}$. Neither value $k\in\{\sigma_{i-1}^{+},\sigma_{i}^{-}\}$ satisfied the check at Line 8, so $\hat{\mu}_{\beta_{i}}(\sigma_{i-1}^{+})<e^{-\nu/2}w_{i}$ and $\hat{\mu}_{\beta_{i-1}}(\sigma_{i}^{-})<e^{-\nu/2}w_{i-1}$. In turn, since Line 6 well-estimates these values, we have $\mu_{\beta_{i}}(\sigma_{i-1}^{+})<w_{i}$ and $\mu_{\beta_{i-1}}(\sigma_{i}^{-})<w_{i-1}$. But now $\mu_{\beta_{i}}(\sigma_{i-1}^{+})\mu_{\beta_{i-1}}(\sigma_{i}^{-})<w_{i-1}w_{i}\leq\mu_{\beta_{i}}(\sigma_{i}^{-})\mu_{\beta_{i-1}}(\sigma_{i-1}^{+})$. This contradicts the Uncrossing Inequality since $\sigma_{i}^{-}\leq\sigma_{i-1}^{+}$ by (I0). ∎
@@
2       call Algorithm 11 with appropriate constants $\nu,\lambda,\tau$ to compute preschedule ${\cal J}$
@@
By Theorem 39(a), we have ${\tt InvWt}(\mathcal{I})\leq 2\rho(n+1)\cdot e^{\nu}\lambda^{3}/\tau$. The term $e^{\nu}\lambda^{3}/\tau$ gets arbitrarily close to $2$ for constants $\nu,\lambda,\tau$ sufficiently close to $0,1,\frac{1}{2}$ respectively.
Finally, by Proposition 39, each iteration of Algorithm 13 returns a non-proper covering schedule with probability at most $\gamma/4$ (irrespective of the choice of $\mathcal{J}$). Thus, the total probability of returning a non-proper covering schedule over all iterations is at most $\sum_{i=0}^{\infty}(3/8)^{i}\gamma/4=2\gamma/5$.
@@ 7 Combinatorial applications
We will examine three combinatorial problems from graph theory: connected subgraphs, matchings, and independent sets. There is essentially only one piece of information we need about the underlying Gibbs distribution, namely, the value $\Delta_{\min}$. Esentially all other problem-specific details can be ignored. Furthermore, for log-concave distributions, there is a simple “automatic” bound on $\Delta_{\min}$; combined with our algorithm for log-concave distributions, this gives a particularly clean algorithmic result.
@@
Similarly, we have $\mu_{b_{1}}(0)\geq\frac{1}{n+1}$. So $b_{i}\in[b_{1},b_{n}]\subseteq[\beta_{\min},\beta_{\max}]$ and $\Delta_{\min}\geq\tfrac{1}{n+1}$ as claimed. We next turn to the bound on $Q(\beta_{\max})$. We have the lower bound $Z(\beta_{\min})\geq c_{0}$. To upper-bound $Z(\beta_{\max})$, we observe from Eq. (22) and from our hypothesis $\beta_{\max}\geq b_{n}$ that for every $k\leq n$, we have
@@
So $Z(\beta_{\max})=\sum_{i}c_{i}e^{i\beta_{\max}}\leq(n+1)c_{n}e^{n\beta_{\max}}$. This implies $Q(\beta_{\max})\leq\frac{e^{n\beta_{\max}}(n+1)c_{n}}{c_{0}}$. By telescoping products, $\frac{c_{n}}{c_{0}}=\prod_{i=1}^{n}\frac{c_{i}}{c_{i-1}}\leq(\frac{c_{1}}{c_{0}})^{n}$, giving $Q(\beta_{\max})\leq e^{n\beta_{\max}}\cdot(n+1)\cdot(\frac{c_{1}}{c_{0}})^{n}\leq e^{n\Gamma}\cdot(n+1)\cdot e^{n\Gamma}\leq e^{q}$. ∎
@@ 7.1 Counting connected subgraphs
Let $G=(V,E)$ be a connected graph. The problem of network reliability is to sample a connected subgraph $G^{\prime}=(V,E^{\prime})$ with probability proportional to $\prod_{f\in E^{\prime}}(1-p(f))\prod_{f\in E-E^{\prime}}p(f)$, for any weighting function $p:E\rightarrow[0,1]$. This can be interpreted as each edge $f$ “failing” independently with probability $p(f)$, and conditioning on the resulting subgraph remaining connected. Equivalently, if we set $p(f)=\frac{e^{\beta}}{1+e^{\beta}}$ for all edges $f$, it can be interpreted as a Gibbs distribution where $c_{i}$ is the number of connected subgraphs of $G$ with $|E|-i$ edges and with $n=|E|-|V|+1$.
Note that $c_{i}$ is the number of independent sets in the co-graphic matroid. By the result of [1], this sequence $c_{i}$ is log-concave. Anari et al. [2] provides an FPRAS for counting independent sets in arbitrary matroids, which would include connected subgraphs, but do not provide concrete complexity estimates for their algorithm. There have also been a series of exact and approximate sampling algorithms for network reliability via Gibbs distributions [12, 11, 7]. The most recent result is due to Chen, Zhang, & Zou [7] which we summarize as follows:
@@
There is an algorithm to approximately sample from the network reliability distribution for any value of $\beta$, up to total variation distance $\rho$, with expected runtime $O(|E|(e^{\beta}+1)\log(|V|/\rho)\log^{2}|V|)$.
@@
Here $c_{0}=1$ and $c_{1}/c_{0}\leq|E|$ and $c_{n-1}/c_{n}\leq|E|$. So we can apply Theorem 40, setting $\beta_{\max}=\log|E|,\beta_{\min}=-\infty$ and $\Gamma=\log|E|$. The algorithm uses $O(\frac{|E|\log^{2}|E|}{\varepsilon^{2}})$ samples in expectation to achieve failure probability $\gamma=O(1)$. Accordingly, we need to run the approximate sampler of Theorem 41 with $\rho=\operatorname*{poly}(n,1/\varepsilon)$. The total algorithm runtime is then $\tilde{O}(|E|^{3}/\varepsilon^{2})$. ∎
@@ 7.2 Counting matchings in high-degree graphs
Consider a graph $G=(V,E)$ with a perfect matching. For $i=0,\dots,|V|/2$, let $c_{i}$ denote the number of $i$-edge matchings. As originally shown in [13], the sequence $c_{i}$ is log-concave. In [18, 19], Jerrum & Sinclair described an MCMC algorithm to approximately sample from the Gibbs distribution on matchings. We quote their most optimized result as follows:
@@
Here $c_{0}=1$ and $n=|V|/2$ and $c_{1}/c_{0}\leq|E|$, and by assumption $c_{n-1}/c_{n}\leq f$. So we can apply Theorem 40 setting $\beta_{\min}=-\log|E|,\beta_{\max}=\log f$, and $\Gamma\leq\max\{\log|E|,\log f\}$. The algorithm uses $O(n\log(|E|f)\log^{2}n/\varepsilon^{2})$ samples in expectation for failure probability $\gamma=O(1)$.
By Theorem 55, we can take $\rho=\operatorname*{poly}(1/n,f,\varepsilon)$ to ensure that the sampling oracle is sufficiently close to the Gibbs distribution. Then Theorem 42 requires $O(|E||V|^{2}f\operatorname*{polylog}(|V|,f,1/\varepsilon))$ time per sample, and overall our FPRAS has runtime of $\tilde{O}(\frac{|E||V|^{3}f}{\varepsilon^{2}})$.
Note also that, as shown in [18], if $G$ has minimum degree at least $|V|/2$ then $c_{|V|/2}>0$ and $c_{|V|/2-1}/c_{|V|/2}\leq O(|V|^{2})$ and clearly $|E|\leq O(|V|^{2})$. ∎
@@ 7.3 Counting independent sets in bounded-degree graphs
Let $G=(V,E)$ be a graph of of maximum degree $D$. A key problem in statistical physics is to efficiently sample the independent sets of $G$. We can consider the Gibbs distribution on independent sets weighted by their size, i.e. each independent set $I\subseteq V$ is sampled with probability $\lambda^{|I|}=e^{\beta|I|}$ for a given parameter $\lambda=e^{\beta}$, and $c_{k}$ is the number of independent sets of cardinality $k$. Note that the Gibbs distribution is not necessarily log-concave.
As shown in [8], this distribution has a critical hardness threshold defined by $\lambda_{c}=\frac{(D-1)^{D-1}}{(D-2)^{D}}\approx e/D$. It is intractable to sample from the Gibbs distribution beyond $\beta>\log\lambda_{c}$ and there is a polynomial-time sampler for the Gibbs distribution for $\beta<\log\lambda_{c}$. We quote the following result:
@@
Let $D\geq 3$ and $\xi>0$ be any fixed constants. There is an algorithm to approximately sample from the Gibbs distribution at $\beta\in[-\infty,\log(\lambda_{c})-\xi]$, up to total variation distance $\rho$, with runtime $O(|V|\log|V|\log\tfrac{1}{\rho})$. [^4] Concretely, [8] describes an MCMC known as Glauber dynamics whose stationary distribution is the Gibbs distribution. They show this Markov chain has mixing time $O(|V|\log|V|)$. For a constant-degree graph, each step of the Glauber dynamics can be executed in $O(1)$ time.
The related problem of estimating counts was considered in [9, 17]. They identified a related threshold value $\alpha_{c}=\frac{\lambda_{c}}{1+(D+1)\lambda_{c}}$ such that for $k>\alpha_{c}|V|$, it is intractable to estimate $c_{k}$ or sample approximately uniformly from size-$k$ independent sets, while on the other hand, for any constants $D\geq 3,\xi>0$ and any $k<(\alpha_{c}-\xi)|V|$, there is a polynomial-time algorithm for both tasks. A key technical result shown in [17] was that, in the latter regime, $\mu_{\beta}$ could be approximated by a normal distribution, i.e. it obeyed a type of Central Limit Theorem.
@@
Let $D\geq 3$ and $\xi>0$ be any fixed constants. There is a constant $\xi^{\prime}>0$ such that, for any $k\leq(\alpha_{c}-\xi)|V|$, there is a value $\beta<\log(\lambda_{c})-\xi^{\prime}$ with $\mu_{\beta}(k)\geq\Omega(1/\sqrt{|V|})$.
This bound alone is all we need for our counting algorithm, which we summarize as follows:
@@
Here $c_{0}=1$ and $n=|V|$. We set $\beta_{\min}=-\infty$ and $\beta_{\max}=\log(\lambda_{c})-\xi^{\prime}$, so $\frac{Q(\beta_{\max})}{Q(\beta_{\min})}\leq\frac{2^{n}e^{\beta_{\max}n}}{c_{0}}$; in particular, since $\beta_{\max}=O(1)$ (for fixed $D$), we can take $q=\Theta(n)$. It suffices to solve ${P^{\delta,\varepsilon}_{\tt count}}$ for $\delta=\Omega(1/\sqrt{n})$. We use the continuous-setting algorithm of Theorem 22; for $\gamma=O(1)$, it has expected sample complexity $O\bigl{(}\frac{q\log n+(\sqrt{q\log n}/\delta)\log\tfrac{q}{\delta}}{\varepsilon^{2}}\bigr{)}=O\bigl{(}\frac{n\log^{3/2}n}{\varepsilon^{2}}\bigr{)}$. The sampler of Theorem 43 then should set $\rho=\operatorname*{poly}(n,1/\varepsilon)$, and the overall runtime is $O\bigl{(}\frac{n^{2}\log^{5/2}n\log(n/\varepsilon)}{\varepsilon^{2}}\bigr{)}=\tilde{O}(n^{2}/\varepsilon^{2}).$

### v11

@@ 6.2 Converting the preschedule into a covering schedule
From Proposition 38, we have ${\tt InvWt}(\mathcal{I})=\sum_{i=0}^{t}\frac{1}{w_{i}/\lambda}\leq\frac{2\lambda(n+1)}{\phi}$. Similarly, the algorithm cost from Line 6 is $\sum_{i=0}^{t}O(w_{i}\log\tfrac{t}{\gamma})\leq O(\frac{2(n+1)}{\phi}\cdot\log\tfrac{n}{\gamma})\leq O(n\rho\log\tfrac{n}{\gamma})$, thus showing part (d).
To show $\mathcal{I}$ is a covering schedule, we need to show $k_{1}<\dots<k_{t}$ and $\beta_{0}=\beta_{\min}<\dots<\beta_{t}=\beta_{\max}$. The condition $\beta_{0}=\beta_{\min},\beta_{t}=\beta_{\max}$ follows from (I1) while the bound $\beta_{i}<\beta_{i+1}$ follows from the loop at Line 1. To show $k_{i}<k_{i+1}$ for $i=1,\dots,t-1$, observe that from (I0) we have $\sigma_{i-1}^{+}\geq\sigma_{i}^{-}$ and $\sigma_{i}^{+}\geq\sigma_{i+1}^{-}$. Due to the loop at Line 3, we must have $k_{i+1}\geq\sigma_{i+1}^{-}>\sigma_{i-1}^{+}\geq k_{i}$ (else segment $\theta_{i}$ would have been discarded).
With probability at least $1-\gamma$, Line 6 well-estimates every value $\sigma_{i}^{+},\sigma_{i}^{-},\sigma_{i+1}^{-},\sigma_{i+1}^{+}$. We claim that, in this case, the algorithm outputs either a proper covering schedule or ERROR, and that the latter case only holds if $\mathcal{J}^{\prime}$, and hence $\mathcal{J}$, is improper.
First, if the algorithm reaches Line 11, then $\hat{\mu}_{\beta_{i}}(k_{i})\geq w_{i}/\sqrt{\lambda}$ and $\hat{\mu}_{\beta_{i-1}}(k_{i})\geq w_{i-1}/\sqrt{\lambda}$ for each $i$. Since Line 6 well-estimates $k_{i}$, this implies $\mu_{\beta_{i}}(k_{i})\geq w_{i}/\lambda$ and $\mu_{\beta_{i-1}}(k_{i})\geq w_{i-1}/\lambda$. So $\mathcal{I}$ is proper.
Next, suppose $\mathcal{J}^{\prime}$ is $\phi$-proper but the algorithm outputs ERROR at an iteration $i$. By definition of $\phi$-properness, we have $\mu_{\beta_{i-1}}(\sigma_{i-1}^{+})\geq w_{i-1}$ and $\mu_{\beta_{i}}(\sigma_{i}^{-})\geq w_{i}$. Since Line 6 well-estimates $\sigma_{i-1}^{+}$ and $\sigma_{i}^{+}$, this implies $\hat{\mu}_{\beta_{i-1}}(\sigma_{i-1}^{+})\geq w_{i-1}/\sqrt{\lambda}$ and $\hat{\mu}_{\beta_{i}}(\sigma_{i}^{-})\geq w_{i}/\sqrt{\lambda}$. Neither value $k\in\{\sigma_{i-1}^{+},\sigma_{i}^{-}\}$ satisfied the check at Line 8, so $\hat{\mu}_{\beta_{i}}(\sigma_{i-1}^{+})<w_{i}/\sqrt{\lambda}$ and $\hat{\mu}_{\beta_{i-1}}(\sigma_{i}^{-})<w_{i-1}/\sqrt{\lambda}$. In turn, since Line 6 well-estimates these values, we have $\mu_{\beta_{i}}(\sigma_{i-1}^{+})<w_{i}$ and $\mu_{\beta_{i-1}}(\sigma_{i}^{-})<w_{i-1}$. But now $\mu_{\beta_{i}}(\sigma_{i-1}^{+})\mu_{\beta_{i-1}}(\sigma_{i}^{-})<w_{i-1}w_{i}\leq\mu_{\beta_{i}}(\sigma_{i}^{-})\mu_{\beta_{i-1}}(\sigma_{i-1}^{+})$. This contradicts the Uncrossing Inequality since $\sigma_{i}^{-}\leq\sigma_{i-1}^{+}$ by (I0). ∎
@@
2       call Algorithm 10 with appropriate constants $\lambda,\tau$ to compute preschedule ${\cal J}$
@@
By Theorem 39(a), we have ${\tt InvWt}(\mathcal{I})\leq 2\rho(n+1)\cdot\lambda^{4}/\tau$. The term $\lambda^{4}/\tau$ gets arbitrarily close to $2$ for constants $\lambda,\tau$ sufficiently close to $1,\frac{1}{2}$ respectively.
By Proposition 39, each iteration of Algorithm 12 returns a non-proper covering schedule with probability at most $\gamma/4$ (irrespective of the choice of $\mathcal{J}$). So the total probability of returning a non-proper covering schedule over all iterations is at most $\sum_{i=0}^{\infty}(3/8)^{i}\gamma/4=2\gamma/5$.
@@ 7 Combinatorial applications
We will examine three problems from graph theory: connected subgraphs, matchings, and independent sets. There is essentially only one piece of information we need about the underlying Gibbs distribution, namely, the value $\Delta_{\min}$; most other problem-specific details can be ignored. Furthermore, log-concave distributions have a simple “automatic” bound on $\Delta_{\min}$; combined with our algorithm for log-concave distributions, this gives a particularly clean algorithmic result.
@@
Similarly, we have $\mu_{b_{1}}(0)\geq\frac{1}{n+1}$. So $b_{i}\in[b_{1},b_{n}]\subseteq[\beta_{\min},\beta_{\max}]$ and $\Delta_{\min}\geq\tfrac{1}{n+1}$ as claimed. We next turn to the bound on $Q(\beta_{\max})$. We have the lower bound $Z(\beta_{\min})\geq c_{0}$. To upper-bound $Z(\beta_{\max})$, observe from Eq. (22) and from our hypothesis $\beta_{\max}\geq b_{n}$ that for every $k\leq n$, that
@@
So $Z(\beta_{\max})=\sum_{i}c_{i}e^{i\beta_{\max}}\leq(n+1)c_{n}e^{n\beta_{\max}}$ and hence $Q(\beta_{\max})\leq\frac{e^{n\beta_{\max}}(n+1)c_{n}}{c_{0}}$. By telescoping products, $\frac{c_{n}}{c_{0}}=\prod_{i=1}^{n}\frac{c_{i}}{c_{i-1}}\leq(\frac{c_{1}}{c_{0}})^{n}$, giving $Q(\beta_{\max})\leq e^{n\beta_{\max}}\cdot(n+1)\cdot(\frac{c_{1}}{c_{0}})^{n}\leq e^{n\Gamma}\cdot(n+1)\cdot e^{n\Gamma}\leq e^{q}$. ∎
@@ 7.1 Counting connected subgraphs
Let $G=(V,E)$ be a connected graph. The problem of network reliability is to sample a connected subgraph $G^{\prime}=(V,E^{\prime})$ with probability proportional to $\prod_{f\in E^{\prime}}\frac{1-p(f)}{p(f)}$, for any weighting function $p:E\rightarrow[0,1]$. This can be interpreted as each edge $f$ “failing” independently with probability $p(f)$, and conditioning on the resulting subgraph remaining connected. Equivalently, if we set $p(f)=\frac{e^{\beta}}{1+e^{\beta}}$ for all edges $f$, it can be interpreted as a Gibbs distribution where $c_{i}$ is the number of connected subgraphs of $G$ with $|E|-i$ edges and with $n=|E|-|V|+1$.
Note that $c_{i}$ is the number of independent sets in the co-graphic matroid. By the result of [1], this sequence $c_{i}$ is log-concave. Anari et al. [2] provides an FPRAS for counting independent sets in arbitrary matroids, which would include connected subgraphs, but do not provide concrete complexity estimates for their algorithm. There have also been a series of exact and approximate sampling algorithms for network reliability via Gibbs distributions [12, 11, 7]. The most recent result is due to [7], which applies to general matroids satisfying certain computational properties. For graph reliability, it provides the following algorithmic result:
@@
There is an algorithm to approximately sample from the Gibbs distribution at $\beta$, up to total variation distance $\rho$, with expected runtime $\tilde{O}(|E|(e^{\beta}+1)\log\tfrac{1}{\rho})$.
@@
Here $c_{0}=1$ and $c_{1}/c_{0}\leq|E|$ and $c_{n-1}/c_{n}\leq|E|$. So we can apply Theorem 40, setting $\beta_{\max}=\log|E|,\beta_{\min}=-\infty$ and $\Gamma=\log|E|$. The algorithm uses $O(\frac{|E|\log^{2}|E|}{\varepsilon^{2}})$ samples in expectation to achieve failure probability $\gamma=O(1)$. Accordingly, we need to run the approximate sampler of Theorem 41 with $\rho=\operatorname*{poly}(1/n,\varepsilon)$. The total algorithm runtime is then $\tilde{O}(|E|^{3}/\varepsilon^{2})$. ∎
@@ 7.2 Counting matchings in high-degree graphs
Consider a graph $G=(V,E)$ with a perfect matching. For $i=0,\dots,v=|V|/2$, let $c_{i}$ denote the number of $i$-edge matchings. As originally shown in [13], the sequence $c_{i}$ is log-concave. In [18, 19], Jerrum & Sinclair described an MCMC algorithm to approximately sample from the Gibbs distribution on matchings. We quote their most optimized result as follows:
@@
Here $c_{0}=1$ and $n=v$ and $c_{1}/c_{0}\leq|E|$, and by assumption $c_{n-1}/c_{n}\leq f$. We apply Theorem 40 setting $\beta_{\min}=-\log|E|,\beta_{\max}=\log f$, and $\Gamma\leq\max\{\log|E|,\log f\}$. The algorithm uses $O(n\log(|E|f)\log^{2}n/\varepsilon^{2})$ samples in expectation for failure probability $\gamma=O(1)$. Accordingly, we need to take the approximate sampler of Theorem 42 with $\rho=\operatorname*{poly}(1/n,\varepsilon,1/f)$. The total algorithm runtime is then $\tilde{O}(\frac{|E||V|^{3}f}{\varepsilon^{2}})$. As shown in [18], if $G$ has minimum degree at least $|V|/2$ then $c_{v}>0$ and $c_{v-1}/c_{v}\leq O(|V|^{2})$ and clearly $|E|\leq O(|V|^{2})$. ∎
@@ 7.3 Counting independent sets in bounded-degree graphs
Let $G=(V,E)$ be a graph of of maximum degree $D$. A key problem in statistical physics is to efficiently sample the independent sets of $G$. We consider the Gibbs distribution on independent sets weighted by their size, i.e. each independent set $I\subseteq V$ is sampled with probability $e^{\beta|I|}$, and $c_{k}$ is the number of cardinality-$k$ independent sets. Note that this distribution is not necessarily log-concave.
As shown in [8], there is a critical hardness threshold defined by $\beta_{c}=\log\bigl{(}\frac{(D-1)^{D-1}}{(D-2)^{D}}\bigr{)}$: it is intractable to sample from the Gibbs distribution beyond $\beta>\beta_{c}$ and there is a polynomial-time sampler for the Gibbs distribution for $\beta<\beta_{c}$. We quote the following result:
@@
Let $D\geq 3$ and $\xi>0$ be any fixed constants. There is an algorithm to approximately sample from the Gibbs distribution at $\beta\in[-\infty,\beta_{c}-\xi]$, up to total variation distance $\rho$, with runtime $O(|V|\log|V|\log\tfrac{1}{\rho})$. [^4] Concretely, [8] describes an MCMC known as Glauber dynamics whose stationary distribution is the Gibbs distribution. They show this Markov chain has mixing time $O(|V|\log|V|)$. For a constant-degree graph, each step of the Glauber dynamics can be executed in $O(1)$ time.
The related problem of estimating counts was considered in [9, 17]. They identified a related threshold value $\alpha_{c}=\frac{e^{\beta_{c}}}{1+(D+1)e^{\beta_{c}}}$ such that for $k>\alpha_{c}|V|$, it is intractable to estimate $c_{k}$ or sample approximately uniformly from size-$k$ independent sets, while on the other hand, for any constants $D\geq 3,\xi>0$ and any $k<(\alpha_{c}-\xi)|V|$, there is a polynomial-time algorithm for both tasks. A key technical result shown in [17] was that, in the latter regime, $\mu_{\beta}$ could be approximated by a normal distribution, i.e. it obeyed a type of Central Limit Theorem.
@@
Let $D\geq 3$ and $\xi>0$ be any fixed constants. There is a constant $\xi^{\prime}>0$ such that, for any $k\leq(\alpha_{c}-\xi)|V|$, there is a value $\beta<\beta_{c}-\xi^{\prime}$ with $\mu_{\beta}(k)\geq\Omega(1/\sqrt{|V|})$.
This bound is all we need for our counting algorithm, which we summarize as follows:
@@
Here $c_{0}=1$ and $n=|V|$. We set $\beta_{\min}=-\infty$ and $\beta_{\max}=\beta_{c}-\xi^{\prime}$. So then $\frac{Z(\beta_{\max})}{Z(\beta_{\min})}\leq\frac{2^{n}e^{\beta_{\max}n}}{c_{0}}$; in particular, since $\beta_{\max}=O(1)$ (for fixed $D$), we can take $q=\Theta(n)$. It suffices to solve ${P^{\delta,\varepsilon}_{\tt count}}$ for $\delta=\Omega(1/\sqrt{n})$. We use the continuous-setting algorithm of Theorem 22; for $\gamma=O(1)$, it has expected sample complexity $O\bigl{(}\frac{q\log n+(\sqrt{q\log n}/\delta)\log\tfrac{q}{\delta}}{\varepsilon^{2}}\bigr{)}=O\bigl{(}\frac{n\log^{3/2}n}{\varepsilon^{2}}\bigr{)}$. The sampler of Theorem 43 then should set $\rho=\operatorname*{poly}(1/n,\varepsilon)$, and the overall runtime is $O\bigl{(}\frac{n^{2}\log^{5/2}n\log(n/\varepsilon)}{\varepsilon^{2}}\bigr{)}=\tilde{O}(n^{2}/\varepsilon^{2}).$